PrettyEyes:

If 10% of x is 20, what is 23% of x? A. 33 B. 46 C. 200 D. 230

Angle:

There are two steps to this problem. The first step is to solve for x. The second step is to answer the question. "10% of x is 20" First, let's convert the 10% into a decimal = 0.10 Then our math problem becomes 0.10 * x = 20

Angle:

What number times 0.1 would equal 20?

PrettyEyes:

um i think 20 times will equal to 0.20

Angle:

hmmm 0.1 times 20 would equal 2 but I want 0.1 times _____ to equal 20

Angle:

Let me try to explain a short cut (but because it is a short cut, it might not make sense). 3 * 5 = 15 for the same reason 15 / 3 = 5 You can agree that both of the things I wrote up there are correct math equations. But using the same format, I can say: 0.10 * x = 20 can be rewritten as x = 20 / 0.10 so now what is 20 divided by 0.1?

PrettyEyes:

um i am so confused

Angle:

what is 20 divided by 0.10?

PrettyEyes:

200 i think

Angle:

awesome! so that means 0.1 * x = 20 then x = 200 just like you said

Angle:

now we want to know 23% x 200

Angle:

23% converted to a decimal is 0.23 what is 0.23 x 200 = ?

PrettyEyes:

46

Angle:

that's your answer! :)